OpenStudy (anonymous):

Simplify using the Distributive property: -7(2x + 1)

OpenStudy (anonymous):

-14x-7

OpenStudy (anonymous):

\[a(b+c) = ab+ac\] \[-7(2x+1)=-7(2x)-7(1) = -14x-7\]